### v1

# HyDeFuse: Provably Convergent Denoiser-Driven Hyperspectral Fusion

Abstract
Hyperspectral (HS) images provide fine spectral resolution but have limited spatial resolution, whereas multispectral (MS) images capture finer spatial details but have fewer bands. HS-MS fusion aims to integrate HS and MS images to generate a single image with improved spatial and spectral resolution. This is commonly formulated as an inverse problem with a linear forward model. However, reconstructing high-quality images using the forward model alone is challenging, necessitating the use of regularization techniques. Over the years, numerous methods have been developed, including wavelets, total variation, low-rank models, and deep neural networks. In this work, we investigate the paradigm of denoiser-driven regularization, where a powerful off-the-shelf denoiser is used for implicit regularization within an iterative algorithm. This has shown much promise but remains relatively underexplored in hyperspectral imaging. Our focus is on a crucial aspect of denoiser-driven algorithm — ensuring convergence of the iterations. It is known that powerful denoisers can produce high-quality reconstructions, but they are also prone to instability and can cause the iterations to diverge. The challenge is to come up with denoisers that come with a convergence guarantee. In this work, we consider a denoiser-driven fusion algorithm, HyDeFuse, which leverages a class of pseudo-linear denoisers for implicit regularization. We demonstrate how the contraction mapping theorem can be applied to establish global linear convergence of HyDeFuse. Additionally, we introduce enhancements to the denoiser that significantly improve the performance of HyDeFuse, making it competitive with state-of-the-art techniques. We validate our theoretical results and present fusion results on publicly available datasets to demonstrate the performance of HyDeFuse.
1 Introduction
Hyperspectral cameras capture images of a scene across a continuous range of spectral bands. The spectral information is useful in applications such as navigation, surveillance, object identification, and medical diagnosis [1]. However, the energy available per band is significantly reduced because the incoming light is divided into numerous narrow bands. This creates an inherent tradeoff between spatial and spectral resolution. Hyperspectral (HS) images generally provide high spectral resolution (i.e., have many bands) but suffer from low spatial resolution. In contrast, multispectral (MS) images capture high spatial resolution but contain limited spectral information. HS-MS fusion techniques aim to combine HS and MS images of the same scene to produce an enhanced image with improved spatial and spectral resolution [2, 3]. The fused image is particularly useful in applications such as clustering [4] and classification [5]. HS-MS fusion has several practical challenges, including image misregistration, atmospheric variations, and inconsistent illumination [6]. Most studies simplify the problem by assuming perfect registration between HS and MS images and consistent atmospheric and illumination conditions. However, even under these idealized settings, obtaining high-quality fused images necessitates using sophisticated reconstruction algorithms.
1.1 Literature review
HS-MS fusion techniques can be categorized into five groups: pansharpening [3, 7], Bayesian [8, 9, 10, 11, 12], multiresolution analysis [13], tensor factorization [14, 15, 16, 17, 18, 19], and deep learning [20, 21, 22, 2, 23]. Of these, Bayesian, tensor factorization, and deep learning have garnered attention due to their state-of-the-art performance. In Bayesian methods, the fusion problem is formulated as a maximum a posteriori (MAP) optimization, incorporating appropriate priors on the ground truth. Commonly used priors include Gaussian priors [24], Markov Random Fields [25], transform-domain sparsity priors [12, 19, 8], and low-rank and nonnegativity priors [26]. On the other hand, deep learning methods attempt to learn the functional relationship between the ground truth and the observed HS-MS images using end-to-end deep networks. They can achieve remarkable results, leveraging architectures such as transformers [27, 28, 29, 30, 31], CNNs [32, 33, 34], and diffusion models [35, 36, 37, 38]. Unfortunately, deep learning relies on large training datasets, which can be difficult to obtain when high-resolution ground truth images are unavailable [2]. Training deep models is computationally intensive, and models trained for specific forward models often lack adaptability [2].
Figure 1: An illustration of instability in denoiser-driven fusion on the Pavia dataset. Specifically, we compare the performance of plug-and-play (PnP) fusion using DnCNN [39] and the proposed CasKD denoiser. The plots track the evolution of PSNR and the distance between successive iterates over iterations. The inset images show the fused results at iterations 1, 100, and 1500. Notably, with DnCNN, the PSNR initially improves but then drops sharply with iterations, ultimately leading to a poor reconstruction. This instability is also highlighted in the second plot, showing the divergence of the iterates $\mathbf{X}_{k}$. This phenomenon is, in fact, peculiar to deep denoisers and has been previously observed for other inverse problems [40]. In contrast, the PSNR steadily increases with our denoiser, and the successive differences decay to zero, which aligns with the theoretical guarantees established in this work.
1.2 Motivation
Recently, state-of-the-art denoisers have been successfully employed for image regularization, achieving state-of-the-art results [41, 42, 43, 44]. Our work focuses on the Plug-and-Play (PnP) method, where an off-the-shelf denoiser is used as a regularizer inside a classical reconstruction algorithm. The denoiser acts as an implicit prior, guiding the iterative process towards a better reconstruction. Compared to end-to-end methods, Plug-and-Play (PnP) allows the decoupling of the forward model and the regularizer. In an end-to-end approach, the forward model is typically utilized during training but not during inference. In contrast, PnP leverages a pretrained denoiser, allowing it to be applied across different forward models. The flip side is that PnP algorithms are strictly not derived from an optimization framework, and the convergence of the iterates is not guaranteed automatically. It is thus not surprising that the convergence of PnP methods has been an active area of research [42, 45, 46, 47, 48, 49]. It is worth noting that the convergence guarantees for deep denoisers often depend on strong technical assumptions that are either impractical or difficult to verify in real-world scenarios.
The choice of the denoiser in PnP can significantly impact the reconstruction quality. Classical denoisers [50, 51, 52, 53, 54, 55, 56] do not require training and offer better interpretability and generalizability. Deep denoisers have recently gained popularity due to their strong denoising capability [39, 57, 58, 53]. In the PnP framework, the denoiser acts more as a regularizer—a smoothing operator—than merely removing “noise”  [59, 46]. Unlike traditional denoising, where the denoiser is applied once to a noisy image, PnP algorithms employ the denoiser iteratively. Consequently, even a relatively weak denoiser can be an effective regularizer, producing a high-quality reconstruction. Rather, the challenge lies in ensuring that the PnP iterations remain stable and do not diverge. As illustrated in Figure 1, powerful deep denoisers such as DnCNN [39] can cause the PnP iterations to become unstable and diverge beyond a point. We have also observed this instability in hyperspectral imaging scenarios, underscoring the need for denoisers that improve image quality and guarantee convergence.
1.3 Contribution
Recent works have demonstrated that a class of pseudo-linear denoisers, known as kernel denoisers, can be used to develop convergent PnP algorithms [52, 49, 42, 60]. The convergence guarantee stems from the unique mathematical properties of kernel denoisers, which are challenging to establish for (black-box) deep denoisers. While kernel denoisers are not as powerful as DnCNN [39] and DRUNET [61], they have been shown to give good reconstructions for inverse problems such as deblurring and superresolution [62, 52, 42, 48]. Kernel denoisers have also been used for convergent hyperspectral fusion [52]. However, as explained in Section 4, the analysis in these works does not apply to our PnP algorithm. This is because we utilize an enhanced denoiser that lacks specific properties (such as symmetry) required in previous analyses. More precisely, our contribution is as follows.
  1. 1.
    Cascaded denoiser. We introduce a hyperspectral denoiser, CasKD, by cascading two existing kernel denoisers [52, 42]. The first denoiser is designed to capture inter-band correlations, leveraging spectral dependencies in hyperspectral images, while the second denoiser exploits in-band correlations. We show that CasKD offers superior denoising and regularization capabilities compared to either denoiser alone, albeit with an increased computational cost. This improvement is significant, as simply cascading two denoisers does not always ensure better performance. We also analyze the mathematical properties of CasKD that are necessary for the convergence analysis.
  2. 2.
    Convergence analysis. We integrate the proposed CasKD denoiser with the proximal gradient descent algorithm [63] to develop a fusion algorithm, HyDeFuse. We prove that HyDeFuse is guaranteed to converge linearly to a unique reconstruction for any arbitrary initialization. To the best of our knowledge, there are not many (iterative) fusion algorithms where the iterates are guaranteed to converge at a geometric rate. Importantly, the analysis allows us to identify the correct step size required to ensure convergence.
  3. 3.
    Validation and comparison. We validate our theoretical results and conduct extensive experiments on publicly available datasets to demonstrate that HyDeFuse is competitive with state-of-the-art fusion methods.
1.4 Organization
We review background materials in Section 2, including the forward model and the loss function for HS-MS fusion. The proposed denoiser and the fusion algorithm are presented in Section 3, where we also discuss its mathematical properties. The core part is Section 4, where we analyze the convergence of HyDeFuse. Finally, in Section 5, we validate our theoretical findings and compare the performance of HyDeFuse with existing fusion algorithms.
2 Background
2.1 Hyperspectral Imaging
The problem of HS-MS fusion can be formulated as an inverse problem with a linear forward model. This formulation is derived from simplifying assumptions about imaging optics, which inherently include nonlinear distortions, sensor noise, and optical aberrations [64, 65]. The sensor plane in hyperspectral imaging consists of an array of detectors that measure the intensity of incoming light at different wavelengths. In MS imaging, each sensor is sensitive to specific regions of the electromagnetic spectrum. In contrast, HS imaging sensors are designed to capture light across several narrow, contiguous spectral bands. A common approach is to model the sensor plane as a two-dimensional lattice, where the sensors are arranged on a sampling grid [65]. In MS imaging, the sensor distribution is denser for a given ground-truth resolution, leading to a higher spatial sampling rate than in HS imaging. As a result, MS images generally exhibit superior spatial resolution. However, each MS pixel integrates information from multiple spectral bands, resulting in spectral degradation. Although HS images capture finer spectral variations over a broader range, their lower sensor density and narrower spectral bands lead to averaging at the sensor plane, making them appear as blurred versions of the high-resolution ground truth. Moreover, due to the lower sampling rate of HS sensors and the reduced energy received per spectral band, spatial downsampling occurs at the sensor level.
2.2 Forward Model
The imaging process described above can be modeled using low-pass filtering and downsampling applied to an (unknown) ground truth image with high spatial and spectral resolution [65, 8]. The inverse problem is to estimate this ground truth from the available HS and MS images. They are naturally represented as 3D tensors, with the three dimensions corresponding to spatial rows, columns, and spectral bands. In the literature, these 3D tensors are often reshaped into 2D matrices for easier handling and analysis, where each column represents the vectorized image of a specific spectral band. The standard forward model is given by
$\mathbf{Y}_{h}=\mathbf{A}\mathbf{Z}+\mathbf{\Theta}_{h}\qquad\mbox{and}\qquad\mathbf{Y}_{m}=\mathbf{Z}\mathbf{R}+\mathbf{\Theta}_{m},$ (1)
where
  • (i)
    $\mathbf{Z}\in\mathbb{R}^{N_{m}\times L_{h}}$ represents the ground truth image with high spectral and spatial resolution, where $N_{m}$ is the number of pixels and $L_{h}$ is the number of spectral bands.
  • (ii)
    $\mathbf{A}=\mathbf{S}\mathbf{B}$ represents spatial degradation, where $\mathbf{S}\in\mathbb{R}^{N_{h}\times N_{m}}$ represents spatial subsampling and $\mathbf{B}\in\mathbb{R}^{N_{m}\times N_{m}}$ represents spatial blurring,
  • (iii)
    $\mathbf{R}\in\mathbb{R}^{L_{h}\times L_{m}}$ represents spectral degradation, with each column representing the spectral response of a specific band,
  • (iv)
    $\mathbf{Y}_{h}\in\mathbb{R}^{N_{h}\times L_{h}}$ and $\mathbf{Y}_{m}\in\mathbb{R}^{N_{m}\times L_{m}}$ are the observed HS and MS images, and
  • (v)
    $\mathbf{\Theta}_{h}\in\mathbb{R}^{N_{h}\times L_{h}}$ and $\mathbf{\Theta}_{m}\in\mathbb{R}^{N_{m}\times L_{m}}$ represent white Gaussian noise.
In the above representation, each column of $\mathbf{Z}$ corresponds to the vectorized form of a spectral band, and the matrices $\mathbf{B},\mathbf{S}$ (resp. $\mathbf{R}$) act on the columns (resp. rows) of $\mathbf{Z}$. They are not explicitly stored as matrices in practice but are applied as operators.
2.3 Variable Reduction
The simplified forward model (1) reduces the fusion problem to solving a system of linear equations. However, note that the number of equations and variables in (1) are $N_{h}L_{h}+N_{m}L_{m}$ and $N_{m}L_{h}$. Typically, $N_{h}\ll N_{m}$ and $L_{m}\ll L_{h}$, meaning the number of unknowns exceeds the number of measurements, making the computation of $\mathbf{Z}$ a highly ill-posed problem [52]. This is precisely where regularization becomes crucial. Moreover, following prior work [8, 66], we use dimensionality reduction to decrease the number of variables. Specifically, we assume that the bands in $\mathbf{Z}$ are highly correlated and that $\mathbf{Z}$ lies in a subspace of lower dimension. Mathematically, this can be expressed as a decomposition,
$\mathbf{Z}=\mathbf{X}\mathbf{E},\qquad\mathbf{X}\in\mathbb{R}^{N_{m}\times L_{s}},\ \mathbf{E}\in\mathbb{R}^{L_{s}\times L_{h}},$ (2)
where the rows of $\mathbf{E}$ form the basis of the low-dimensional subspace. In other words, the model (1) becomes
$\mathbf{Y}_{h}=\mathbf{A}\mathbf{X}\mathbf{E}+\mathbf{\Theta}_{h}\qquad\mbox{and}\qquad\mathbf{Y}_{m}=\mathbf{X}\mathbf{E}\mathbf{R}+\mathbf{\Theta}_{m},$ (3)
where $\mathbf{X}$ is the unknown (or latent) variable. The reconstructed image is given by $\widehat{\mathbf{Z}}=\widehat{\mathbf{X}}\mathbf{E}$, where $\widehat{\mathbf{X}}$ is obtained by solving (3).
A standard method for estimating $\mathbf{E}$ from the observed image $\mathbf{Y}_{h}$ is using the singular value decomposition [43, 8]. Specifically, we first upsample $\mathbf{Y}_{h}$ to the spatial resolution of $\mathbf{Z}$. The resulting image, say $\mathbf{Y}$, is a surrogate of the ground truth. We perform its singular value decomposition
$\mathbf{Y}=\sum_{j=1}^{r}\sigma_{j}\boldsymbol{u}_{j}\boldsymbol{v}_{j}^{\top},$
where $\sigma_{1}\geqslant\cdots\geqslant\sigma_{r}$ are its singular values, and $\boldsymbol{u}_{1},\ldots,\boldsymbol{u}_{r}\in\mathbb{R}^{N_{m}}$ and $\boldsymbol{v}_{1},\ldots,\boldsymbol{v}_{r}\in\mathbb{R}^{L_{h}}$ are the left and right singular vectors, and $r$ is the rank of $\mathbf{Y}$. We choose some $L_{s}\leqslant r$ as the dimension of the targeted subspace and define
$\mathbf{E}=\Big{[}\boldsymbol{v}_{1}\cdots\boldsymbol{v}_{L_{s}}\Big{]}^{\top}\in\mathbb{R}^{L_{s}\times L_{h}}$
as the basis in (3). We will later use the property that $\mathbf{E}\mathbf{E}^{\top}=\mathbf{I}$.
3 Denoiser-Driven Fusion
3.1 Definitions and Notations
Generally, we use bold lowercase letters ($\boldsymbol{x},\boldsymbol{e}$, etc.) to represent vectors, bold uppercase letters ($\mathbf{X},\mathbf{Z}$, etc.) to represent matrices, and calligraphic letters ($\mathcal{T},\mathcal{W}$, etc.) to denote operators acting on matrices. For any $\boldsymbol{x}\in\mathbb{R}^{n}$, $\operatorname*{diag}(\boldsymbol{x})\in\mathbb{R}^{n\times n}$ denotes a diagonal matrix with $\boldsymbol{x}$ as its diagonal entries. The identity matrix is denoted by $\mathbf{I}$, where the size should be clear from the context. We say that $\mathbf{X}$ is a fixed point of $\mathcal{T}$ if $\mathcal{T}(\mathbf{X})=\mathbf{X}$; the set of fixed points of $\mathcal{T}$ is denoted as $\mathrm{fix}\,(\mathcal{T})$. If $\mathcal{T}$ is linear, $\mathrm{fix}\,(\mathcal{T})$ corresponds to the eigenspace associated with the eigenvalue $1$.
We use $\mathbb{H}=\mathbb{R}^{N_{m}\times L_{s}}$ to denote the space of the matrix-valued variable $\mathbf{X}$ in (3). The reconstruction algorithm will operate in this vector space. To establish convergence, we will work with an inner product in $\mathbb{H}$ and the associated induced norm. In particular, it will be convenient to work with the following inner product on $\mathbb{H}$,
$\big{\langle}\mathbf{X}_{1},\mathbf{X}_{2}\big{\rangle}_{\mathbb{H}}=\mathrm{tr}(\mathbf{X}_{1}^{\top}\mathbf{X}_{2})\qquad(\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}),$ (4)
where $\mathrm{tr}$ is the trace operator. The columns of $\mathbf{X}\in\mathbb{H}$ are the spectral bands, so the inner product in (4) amounts to computing the standard dot product between the corresponding bands of $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$, followed by summation over all bands. The norm induced by this inner product is given by
$\|\mathbf{X}\|_{\mathbb{H}}=\sqrt{\langle\mathbf{X},\mathbf{X}\rangle}_{\mathbb{H}}\qquad(\mathbf{X}\in\mathbb{H}).$
We use $\mathcal{L}(\mathbb{H})$ to denote the class of linear operators that map $\mathbb{H}$ into $\mathbb{H}$. We use $\sigma(\mathbf{M})$ and $\sigma(\mathcal{T})$ for the spectrum (set of eigenvalues) of a matrix $\mathbf{M}$ or an operator $\mathcal{T}$. We use $\sigma_{\max}(\mathbf{M})$ for the largest singular value of $\mathbf{M}$. This should not be confused with the notation for the spectrum; the distinction should be clear from the context. We will use the inequality $\|\mathbf{A}\mathbf{B}\|_{\mathbb{H}}\leqslant\sigma_{\max}(\mathbf{A})\|\mathbf{B}\|_{\mathbb{H}}$ in our analysis.
We will use the concept of a self-adjoint operator, a natural generalization of the concept of a symmetric matrix [67]. An operator $\mathcal{T}$ on $(\mathbb{H},\big{\langle}\cdot,\cdot\rangle_{\mathbb{H}})$ is said to be self-adjoint if
$\big{\langle}\mathcal{T}(\mathbf{X}_{1}),\mathbf{X}_{2}\big{\rangle}_{\mathbb{H}}=\big{\langle}\mathbf{X}_{1},\mathcal{T}(\mathbf{X}_{2})\big{\rangle}_{\mathbb{H}}\qquad(\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}).$
A symmetric matrix or a self-adjoint operator is positive semidefinite if all the eigenvalues are nonnegative. An operator $\mathcal{T}$ on $\mathbb{H}$ is nonexpansive if $\|\mathcal{T}(\mathbf{X}_{1})-\mathcal{T}(\mathbf{X}_{2})\|_{\mathbb{H}}\leqslant\|\mathbf{X}_{1}-\mathbf{X}_{2}\|_{\mathbb{H}}$ for all $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}$. In particular, a matrix $\mathbf{M}$ is nonexpansive if $\sigma_{\max}(\mathbf{M})\leqslant 1$. A function $\ell:\mathbb{H}\to\mathbb{R}$ is said to be convex if $\ell(\theta\mathbf{X}_{1}+(1-\theta)\mathbf{X}_{2})\leqslant\theta\ell(\mathbf{X}_{1})+(1-\theta)\ell(\mathbf{X}_{2})$ for all $\theta\in(0,1)$ and $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}$.
Figure 2: The processing blocks of HyDeFuse. The hyperspectral image $\mathbf{Y}_{h}$ is upscaled (interpolated) to match the spatial dimension of the multispectral image, which is used to estimate the spectral subspace $\mathbf{E}$. The loss function $\ell$ is derived from the observed HS and MS images and the forward operators $\mathbf{A}$ and $\mathbf{R}$. PnP-PGD is an iterative algorithm where each iteration consists of a gradient step on the loss function followed by regularization using denoiser $\mathcal{D}$. The output of PnP-PGD is $\widehat{\mathbf{X}}$ and $\widehat{\mathbf{Z}}=\widehat{\mathbf{X}}\mathbf{E}$ is the output of HyDeFuse.
3.2 Plug-and-Play Fusion
The PnP framework [42] has its origin in the standard variational framework for solving (3), namely,
$\underset{\mathbf{X}\in\mathbb{H}}{\min}\ \ell(\mathbf{X})+\varphi(\mathbf{X}),$ (5)
where $\ell:\mathbb{H}\to\mathbb{R}$ is the model-based loss function,
$\ell(\mathbf{X})=\frac{1}{2}\|\mathbf{A}\mathbf{X}\mathbf{E}-\mathbf{Y}_{h}\|_{\mathbb{H}}^{2}+\frac{\lambda}{2}\|\mathbf{X}\mathbf{E}\mathbf{R}-\mathbf{Y}_{m}\|_{\mathbb{H}}^{2},$ (6)
and $\varphi:\mathbb{H}\to\mathbb{R}$ is a (convex) regularizer. Since the regularizer $\varphi$ is typically non-smooth, proximal algorithms are used to perform the optimization [65, 8, 68, 69]. For instance, since the loss is differentiable, we can use the proximal gradient descent (PGD) algorithm. Given an initial estimate $\mathbf{X}_{0}\in\mathbb{H}$, the PGD iterations are given by
$\mathbf{X}_{k+1}=\mathrm{prox}_{\gamma\varphi}\big{(}\mathbf{X}_{k}-\gamma\nabla\!\ell(\mathbf{X}_{k})\big{)}\qquad(k=0,1,\ldots),$ (7)
where $\gamma>0$ is the step size, $\nabla\!\ell$ is the gradient of $\ell$, and $\mathrm{prox}_{\gamma\varphi}:\mathbb{H}\to\mathbb{H}$ is the proximal operator of $\gamma\varphi$,
$\mathrm{prox}_{\gamma\varphi}(\mathbf{X})=\underset{\mathbf{H}\in\mathbb{H}}{\mbox{argmin}}\ \frac{1}{2}\|\mathbf{H}-\mathbf{X}\|_{\mathbb{H}}^{2}+\gamma\varphi(\mathbf{H}).$ (8)
Instead of relying on a classical regularizer and its proximal operator that acts as a Gaussian denoiser [42], the ingenious idea in PnP was to replace the proximal operator in (7) with an arbitrary denoising operator $\mathcal{D}:\mathbb{H}\to\mathbb{H}$. That is, the update in PnP is performed as
$\mathbf{X}_{k+1}=\mathcal{D}\big{(}\mathbf{X}_{k}-\gamma\nabla\!\ell(\mathbf{X}_{k})\big{)}\qquad(k=0,1,\ldots).$ (9)
The point is that denoisers such as BM3D [70], DnCNN [39, 43], and DRUNet [61] are more potent than the proximal operator (8) corresponding to traditional regularizers such as total variation, wavelets, and low-rank [8]. The hope is that a more potent denoiser $\mathcal{D}$ in (9) can produce better reconstructions. However, the challenge lies in the fact that the transition from (7) to (9) is carried out in an ad-hoc manner, with no guarantee that the sequence of iterates $\{\mathbf{X}_{k}\}$ generated by (9) will converge to a meaningful reconstruction. The example in Figure 1 shows that the iterative process can actually fail, leading to divergence. Thus, we need to work with a denoiser $\mathcal{D}$ that not only gives good reconstructions but for which the iterates $\{\mathbf{X}_{k}\}$ are always convergent.
In the rest of the paper, we will refer to (9) as PnP-PGD. We note that this is not a new algorithm and has been used in prior works on PnP [59, 49]. The novelty of the present work is the denoiser we propose next and the proof of convergence of PnP-PGD for this particular denoiser.
We note that the gradient $\nabla\!\ell$ in (7) is with respect to the inner product (4), i.e., the matrix $\nabla\!\ell(\mathbf{X})$ is such that
$\ell(\mathbf{X}+\mathbf{H})=\ell(\mathbf{X})+\langle\nabla\!\ell(\mathbf{X}),\mathbf{H}\big{\rangle}_{\mathbb{H}}+r(\mathbf{H}),$ (10)
where the residual $r:\mathbb{H}\to\mathbb{R}$ is such that $r(\mathbf{H})/\|\mathbf{H}\|_{\mathbb{H}}\to 0$ as $\|\mathbf{H}\|_{\mathbb{H}}\to 0$. Substituting (6) into (10) and performing some calculations, we can verify that
$\nabla\!\ell(\mathbf{X})=\mathbf{A}^{\top}\!\mathbf{A}\mathbf{X}\mathbf{E}\mathbf{E}^{\top}-\mathbf{A}^{\top}\mathbf{Y}_{h}\mathbf{E}^{\top}+\lambda\mathbf{X}(\mathbf{E}\mathbf{R})(\mathbf{E}\mathbf{R})^{\top}-\lambda\mathbf{Y}_{m}(\mathbf{E}\mathbf{R})^{\top}.$
Moreover, since $\mathbf{E}\mathbf{E}^{\top}=\mathbf{I}$, we have
$\nabla\!\ell(\mathbf{X})=\mathbf{A}^{\top}\!\mathbf{A}\mathbf{X}-\mathbf{A}^{\top}\mathbf{Y}_{h}\mathbf{E}^{\top}+\lambda\mathbf{X}(\mathbf{E}\mathbf{R})(\mathbf{E}\mathbf{R})^{\top}-\lambda\mathbf{Y}_{m}(\mathbf{E}\mathbf{R})^{\top}.$ (11)
3.3 Kernel Denoiser
We now describe the denoiser $\mathcal{D}$ in (9). The noise in hyperspectral images is assumed to be uncorrelated both within and across bands [65], but the bands exhibit strong correlations, with numerous similar patches present within each band. To exploit the correlation across bands, we first denoise the image using a kernel denoiser proposed in [52]. Furthermore, to exploit the presence of similar patches within each band, we process each band of the denoised image separately using another kernel denoiser [42]. Specifically, if we denote the first denoiser as an operator $\mathcal{W}:\mathbb{H}\to\mathbb{H}$ and the second denoiser as $\mathcal{V}:\mathbb{H}\to\mathbb{H}$, then the proposed denoiser $\mathcal{D}$ is given by
$\mathcal{D}=\mathcal{V}\circ\mathcal{W}.$ (12)
We will refer to $\mathcal{W}$ as the high-dimensional kernel denoiser, $\mathcal{V}$ as the bandwise kernel denoiser, and $\mathcal{D}$ as CasKD (cascaded kernel denoiser).
The properties of $\mathcal{W}$ and $\mathcal{V}$ will play a crucial role in the convergence analysis. Therefore, we first give the mathematical description of these denosiers. Both $\mathcal{W}$ and $\mathcal{V}$ are examples of kernel denoisers [71], where each pixel is denoised using a weighted average of its neighboring pixels. The weights are determined using a kernel (or affinity) function, hence the name “kernel denoiser.” The main difference between $\mathcal{W}$ and $\mathcal{V}$ is in the definition of the kernel function.
We first give the construction of the bandwise denoiser $\mathcal{V}$, which is more straightforward. For each band $1\leqslant b\leqslant L_{s}$, we define a kernel matrix $\mathbf{K}^{(b)}$, whose components are
$\mathbf{K}^{(b)}_{ij}=h(i-j)\phi(\boldsymbol{p}_{i}-\boldsymbol{p}_{j})\qquad(1\leqslant i,j\leqslant N_{m}),$ (13)
where
  • $\phi$ is a multivariate Gaussian (RBF) kernel function,
  • $h$ is a symmetric hat function supported on a square window around the origin, and
  • $\boldsymbol{p}_{i}$ and $\boldsymbol{p}_{j}$ are the (vectorized) patches around pixels $i$ and $j$ extracted from a guide image, e.g., a surrogate of the ground truth constructed from the observed images [59, 60].
By construction, $\mathbf{K}^{(b)}_{ij}$ takes on large values when pixels $i$ and $j$ and the corresponding patches $\boldsymbol{p}_{i}$ and $\boldsymbol{p}_{j}$ are close. A box function could be used instead of a hat function in (13); the latter guarantees that the kernel matrix is positive semidefinite [42]. In the standard construction of a kernel denoiser [71], the weight matrix is obtained by normalizing $\mathbf{K}^{(b)}$,
$\mathbf{D}=\operatorname*{diag}(\mathbf{K}^{(b)}\,\boldsymbol{e}),\quad\mathbf{W}^{(b)}=\mathbf{D}^{-1}\,\mathbf{K}^{(b)}.$
where $\boldsymbol{e}$ is the all-ones vector of appropriate length. The problem is that although $\mathbf{K}^{(b)}$ is symmetric, $\mathbf{W}^{(b)}$ may not remain symmetric after normalization –– the product of two symmetric matrices need not be symmetric. As originally proposed in [42], a symmetric weight matrix can be constructed using the following formula, which we continue to denote by the same symbol:
$\mathbf{W}^{(b)}=\frac{1}{\nu}\mathbf{D}^{-\frac{1}{2}}\mathbf{K}^{(b)}\mathbf{D}^{-\frac{1}{2}}+\operatorname*{diag}\Big{(}\boldsymbol{e}-\frac{1}{\nu}\,\hat{\boldsymbol{e}}\Big{)}$, (14)
where
$\mathbf{D}=\operatorname*{diag}(\mathbf{K}^{(b)}\,\boldsymbol{e}),\ \hat{\boldsymbol{e}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{K}^{(b)}\mathbf{D}^{-\frac{1}{2}}\boldsymbol{e},\ \nu=\max_{i}\ \hat{\boldsymbol{e}}_{i}.$
The bandwise denoiser $\mathcal{V}:\mathbb{H}\to\mathbb{H}$ is given by the linear transform
$\mathcal{V}(\mathbf{X})_{b}=\mathbf{W}^{(b)}\mathbf{X}_{b}\qquad(\mathbf{X}\in\mathbb{H}),$ (15)
where $\mathbf{X}_{b}\in\mathbb{R}^{N_{m}}$ and $\mathcal{V}(\mathbf{X})_{b}$ denote the $b$-th band of the input and output of the denoiser.
The construction of the high-dimensional denoiser $\mathcal{W}$ is similar but more intricate. Unlike (15), a single linear transform is applied uniformly across all bands. The main steps in the construction are as follows (see [52] for more details):
  1. 1.
    Given the input hyperspectral image represented as a matrix $\mathbf{X}$ of size $N_{m}\times L_{s}$, we extract patches around each pixel while considering all the $L_{s}$ bands. Specifically, for each pixel $1\leqslant i\leqslant N_{m}$, we have a patch vector $\boldsymbol{p}_{i}$ of dimension $L_{s}k^{2}$, assuming that a $k\times k$ neighborhood is used around each pixel.
  2. 2.
    To extract the inter and intra-band correlations among the patches $\boldsymbol{p}_{1},\ldots,\boldsymbol{p}_{N_{s}}$, we group them into $C$ clusters with centroids $\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{C}$.
  3. 3.
    The kernel matrix $\mathbf{K}$ of size $N_{m}\times N_{m}$ is defined to be
    $\mathbf{K}_{i,j}=h(i-j)\,\sum_{c=1}^{C}\phi(\boldsymbol{p}_{i}-\boldsymbol{\mu}_{c})\phi(\boldsymbol{p}_{j}-\boldsymbol{\mu}_{c}),\qquad(1\leqslant i,j\leqslant N_{m}),$ (16)
    where $\phi$ is again a Gaussian kernel. Note that $\mathbf{K}_{i,j}$ assumes relatively large values when $\boldsymbol{p}_{i}$ and $\boldsymbol{p}_{j}$ belong to the same cluster.
  4. 4.
    We define a symmetric weight matrix $\mathbf{W}$ of size $N_{m}\times N_{m}$ using (14), i.e., we set
    $\mathbf{W}=\frac{1}{\nu}\mathbf{D}^{-\frac{1}{2}}\mathbf{K}\mathbf{D}^{-\frac{1}{2}}+\operatorname*{diag}\Big{(}\boldsymbol{e}-\frac{1}{\nu}\,\hat{\boldsymbol{e}}\Big{)}$, (17)
    where
    $\mathbf{D}=\operatorname*{diag}(\mathbf{K}\,\boldsymbol{e}),\ \hat{\boldsymbol{e}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{K}\mathbf{D}^{-\frac{1}{2}}\boldsymbol{e},\ \nu=\max_{i}\ \hat{\boldsymbol{e}}_{i}.$
  5. 5.
    The high-dimensional denoiser $\mathcal{W}:\mathbb{H}\to\mathbb{H}$ is given by
    $\mathcal{W}(\mathbf{X})=\mathbf{W}\mathbf{X}\qquad(\mathbf{X}\in\mathbb{H}).$ (18)
    In other words, the same weight matrix $\mathbf{W}$ is applied uniformly across all bands of $\mathbf{X}$.
The primary motivation behind the construction in (16) is that it facilitates the development of a fast convolution-based algorithm for computing (18). Details on this and other implementation aspects can be found in [52]. Rather, what is more important is the following observation.
Proposition 1.
The matrices $\mathbf{W}^{(b)}$ and $\mathbf{W}$ in (14) and (17) are symmetric positive semidefinite, stochastic, and irreducible.
Stochastic means $\mathbf{W}_{ij}$ are nonnegative, and all the rows sum to one. By irreducible, we mean that for all $i,j$, there exists $k\geqslant 1$, such that $(\mathbf{W}^{k})_{ij}>0$.
Proposition 1 is a key result that relies on the intricate construction of the denoisers in Section 3.3. Indeed, starting with a kernel function, constructing a denoiser with the above mathematical properties is a non-trivial task. Later, we will demonstrate how these properties manifest in the corresponding hyperspectral denoisers $\mathcal{V}$ and $\mathcal{W}$.
The technical details behind Proposition 1 can be found in [42, Sec. IV]. However, one can directly verify from (17) that $\mathbf{W}$ is symmetric, nonnegative, and $\mathbf{W}\boldsymbol{e}=\boldsymbol{e}$ (stochastic). It is also not difficult to show that $\mathbf{W}$ –– for that matter, any operator that performs local averaging on the signal using strictly positive weights –– is irreducible. The only tricky property is the positive definite property; this can be deduced using the Fourier transform and Bochner’s theorem [42].
We note that different orderings and combinations of $\mathcal{V}$ and $\mathcal{W}$, such as $\mathcal{W}\circ\mathcal{V}$, could have been used in (12). However, we empirically found that $\mathcal{V}\circ\mathcal{W}$ gives slightly better results. In either case, the denoising performance surpasses that of using $\mathcal{V}$ or $\mathcal{W}$ alone. This is demonstrated in Table 1 and Figure 3. Moreover, since ultimately we wish to use CasKD to perform fusion (and not for standalone denoising), we have compared its regularization capacity for HyDeFuse with the component denoisers in Table 9. We see a significant gap in the fusion quality.
Noise variance Denoiser PSNR (Pavia) UIQI (Pavia) PSNR (Paris) UIQI (Paris)
noisy 22.10 0.68 22.10 0.68
$\mathcal{W}$ 35.45 0.973 34.52 0.961
$\mathcal{W}\circ\mathcal{W}$ 36.05 0.978 34.71 0.971
20/255 $\mathcal{W}\circ\mathcal{V}$ 38.01 0.986 36.87 0.982
$\mathcal{V}$ 37.50 0.984{0.984} 36.80{36.80} 0.981{0.981}
$\mathcal{V}\circ\mathcal{V}$ 37.50 0.984 36.80 0.981
CasKD $\mathbf{38.06}$ $\mathbf{0.987}$ $\mathbf{37.06}$ $\mathbf{0.982}$
noisy 14.14 0.29 14.14 0.28
$\mathcal{W}$ 29.15 0.897 28.36 0.880
$\mathcal{W}\circ\mathcal{W}$ 30.24 0.920 28.75 0.887
50/255 $\mathcal{W}\circ\mathcal{V}$ 32.38 0.955 30.73 0.915
$\mathcal{V}$ $31.53$ $0.942$ $30.40$ $0.922$
$\mathcal{V}\circ\mathcal{V}$ 31.53 0.942 30.40 0.922
CasKD $\mathbf{32.43}$ $\mathbf{0.955}$ $\mathbf{30.86}$ $\mathbf{0.931}$
Table 1: Comparison of denoising performance on Pavia and Paris. We see that CasKD generally does better than its components $\mathcal{V}$ and $\mathcal{W}$ and their other combinations. The top two methods are in bold and the best method is underlined.
(a) ground truth
(b) noisy
(c) $\mathcal{W}$
(d) $\mathcal{V}$
(e) CasKD
Figure 3: Visual comparison of CasKD with its component denoisers $\mathcal{V}$ and $\mathcal{W}$ for Gaussian noise with standard deviation $0.078$.
4 Convergence Analysis
In this section, we establish global linear convergence of (9) with (12) as the denoiser. Our analysis builds on the convergence framework presented in [49]. However, there are important technical differences that prevent the direct application of the results from [49]. These differences are best explained after presenting the technical results, so we postpone this discussion until the end of the section.
4.1 Fixed Point Iteration
The starting point in the analysis is to view (9) as a fixed point iteration. Specifically, we can write (9) as
$\mathbf{X}_{k+1}=\mathcal{P}(\mathbf{X}_{k})+\mathbf{Q}$ (19)
where $\mathcal{P}\in\mathcal{L}(\mathbb{H})$ and $\mathbf{Q}\in\mathbb{H}$ are given by
$\mathcal{P}(\mathbf{X})=\mathcal{D}\big{(}\mathbf{X}-\gamma\big{(}\mathbf{A}^{\top}\!\mathbf{A}\mathbf{X}+\lambda\mathbf{X}(\mathbf{E}\mathbf{R})(\mathbf{E}\mathbf{R})^{\top}\big{)}\big{)},$
and
$\mathbf{Q}=\mathcal{D}\left(\gamma\mathbf{A}^{\top}\mathbf{Y}_{h}\mathbf{E}^{\top}+\gamma\lambda\mathbf{Y}_{m}(\mathbf{E}\mathbf{R})^{\top}\right).$
Thus, we can view (19) as a fixed point iteration
$\mathbf{X}_{k+1}=\mathcal{T}(\mathbf{X}_{k}),\qquad\mathcal{T}(\mathbf{X}):=\mathcal{P}(\mathbf{X})+\mathbf{Q},$ (20)
We can further decompose $\mathcal{P}$ as
$\mathcal{P}=\mathcal{D}\circ\mathcal{G},$ (21)
where $\mathcal{G}\in\mathcal{L}(\mathbb{H})$ is given by
$\mathcal{G}(\mathbf{X})=\mathbf{X}-\gamma\left(\mathbf{A}^{\top}\!\mathbf{A}\mathbf{X}+\lambda\mathbf{X}(\mathbf{E}\mathbf{R})(\mathbf{E}\mathbf{R})^{\top}\right).$ (22)
We will refer to $\mathcal{G}$ as the gradient step operator and the affine map $\mathcal{T}$ as the fixed point operator.
We will later show that the operators $\mathcal{D}$ and $\mathcal{G}$ are nonexpansive. Consequently, $\mathcal{P}=\mathcal{D}\circ\mathcal{G}$, and hence $\mathcal{T}$, is nonexpansive. However, this is not sufficient to establish convergence of $\{\mathbf{X}_{k}\}$ generated by $\mathbf{X}_{k+1}=\mathcal{T}(\mathbf{X}_{k})$ for any arbitrary $\mathbf{X}_{0}\in\mathbb{H}$. A sufficient condition is that $\mathcal{P}$ (and hence $\mathcal{T}$) is a contraction operator, i.e., there exists $\mu\in[0,1)$ such that
$\|\mathcal{P}(\mathbf{X}_{1})-\mathcal{P}(\mathbf{X}_{2})\|_{\mathbb{H}}\leqslant\mu\,\|\mathbf{X}_{1}-\mathbf{X}_{2}\|_{\mathbb{H}}\qquad(\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}).$ (23)
If $\mathcal{T}$ is a contraction, we use the contraction mapping theorem [72] to conclude that the fixed point iteration $\mathbf{X}_{k+1}=\mathcal{T}(\mathbf{X}_{k})$ converges to a unique fixed point for any initialization $\mathbf{X}_{0}\in\mathbb{H}$. The smallest possible $\mu$ in (23) is called the contraction factor of $\mathcal{T}$.
The rest of this section is dedicated to proving that $\mathcal{T}$ is a contraction operator. As noted above, it suffices to show that $\mathcal{P}$ is a contraction. Moreover, since $\mathcal{P}$ is linear, (23) is equivalent to
$\|\mathcal{P}(\mathbf{X})\|_{\mathbb{H}}\leqslant\mu\,\|\mathbf{X}\|_{\mathbb{H}}\qquad(\mathbf{X}\in\mathbb{H}).$ (24)
We have the following simple observation in this regard.
Proposition 2.
Suppose $\mathbb{H}$ is a finite-dimensional vector space and $\mathcal{P}\in\mathcal{L}(\mathbb{H})$. Then $\mathcal{P}$ is a contraction if and only if $\|\mathcal{P}(\mathbf{X})\|<\|\mathbf{X}\|$ for all nonzero $\mathbf{X}\in\mathbb{H}$. In particular, the contraction factor $\mu$ in (23) is given by
$\mu=\max\Big{\{}\|\mathcal{P}(\mathbf{X})\|_{\mathbb{H}}:\ \mathbf{X}\in\mathbb{H},\,\|\mathbf{X}\|_{\mathbb{H}}=1\Big{\}}.$ (25)
We remark that this need not hold if $\mathcal{P}$ is not linear. One direction of Proposition 2 is obvious. In the other direction, it suffices to show $\mu<1$. Now, the set of $\mathbf{X}\in\mathbb{H}$ such that $\|\mathbf{X}\|_{\mathbb{H}}=1$ forms a compact subset of $\mathbb{H}$ and the map $\mathbf{X}\mapsto\|\mathcal{P}(\mathbf{X})\|_{\mathbb{H}}$ is continuous. Therefore, by the extreme value theorem, there exists $\mathbf{X}^{*}\in\mathbb{H}$ such that $\|\mathbf{X}^{*}\|_{\mathbb{H}}=1$ and $\mu=\|\mathcal{P}(\mathbf{X}^{*})\|_{\mathbb{H}}<\|\mathbf{X}^{*}\|_{\mathbb{H}}=1$.
4.2 Constituent Operators
To show that $\mathcal{P}$ is a contraction, we will establish some special properties of the underlying operators $\mathcal{D}$ and $\mathcal{G}$. Note that we can write (22) as
$\mathcal{G}(\mathbf{X})=\mathbf{X}-\gamma\mathcal{K}(\mathbf{X}),\quad\mathcal{K}(\mathbf{X}):=\mathbf{P}_{1}\mathbf{X}+\mathbf{X}\mathbf{P}_{2},$ (26)
where
$\mathbf{P}_{1}=\mathbf{A}^{\top}\!\mathbf{A}\quad\mbox{and}\quad\mathbf{P}_{2}=\lambda(\mathbf{E}\mathbf{R})(\mathbf{E}\mathbf{R})^{\top}.$ (27)
We claim that $\mathcal{K}$ is self-adjoint on $\mathbb{H}$. Indeed, for any $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}$,
$\big{\langle}\mathcal{K}(\mathbf{X}_{1}),\mathbf{X}_{2}\big{\rangle}_{\mathbb{H}}$ $=$ $\mathrm{tr}\left((\mathbf{P}_{1}\mathbf{X}_{1}+\mathbf{X}_{1}\mathbf{P}_{2})^{\top}\mathbf{X}_{2}\right)$
$=$ $\mathrm{tr}(\mathbf{X}_{1}^{\top}\mathbf{P}_{1}\mathbf{X}_{2})+\mathrm{tr}(\mathbf{P}_{2}\mathbf{X}_{1}^{\top}\mathbf{X}_{2})$
$=$ $\mathrm{tr}(\mathbf{X}_{1}^{\top}\mathbf{P}_{1}\mathbf{X}_{2})+\mathrm{tr}(\mathbf{X}_{1}^{\top}\mathbf{X}_{2}\mathbf{P}_{2})$
$=$ $\big{\langle}\mathbf{X}_{1},\mathcal{K}(\mathbf{X}_{2})\big{\rangle}_{\mathbb{H}},$
where we have used the cyclic property of trace and the fact that $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are symmetric matrices.
Since $\mathcal{K}$ is self-adjoint, it follows from (26) that the gradient step operator $\mathcal{G}$ is self-adjoint. This implies that the eigenvalues of $\mathcal{G}$ are real. We will next show that if the step size $\gamma$ in (22) is sufficiently small, namely if
$0<\gamma<2/\beta,\quad\beta:=\sigma_{\max}(\mathbf{A})^{2}+\lambda\sigma_{\max}(\mathbf{E}\mathbf{R})^{2}.$ (28)
then $\mathcal{G}$ is nonexpansive. Since $\mathcal{G}$ is self-adjoint, this would imply that $\sigma(\mathcal{G})\subset[-1,1]$. Furthermore, we will show that $1\notin\sigma(\mathcal{G})$ if $\gamma$ is in the range given by (28). Specifically, we have the following result.
Proposition 3.
For the step size choice in (28), $\mathcal{G}$ is nonexpansive and $\sigma(\mathcal{G})\subset(-1,1]$.
The nonexpansive of the gradient step operator is a standard result in convex optimization [72]. This relies on the Ballion-Hadded theorem [72], a fundamental result about differentiable convex functions that is applicable for our loss function (6). A function $\ell:\mathbb{H}\to\mathbb{R}$ is said to be $\beta$-smooth if it is differentiable and its gradient $\nabla\!\ell$ is $\beta$-Lipschitz, i.e.,
$\|\nabla\!\ell(\mathbf{X}_{1})-\nabla\!\ell(\mathbf{X}_{2})\|_{\mathbb{H}}\leqslant\beta\|\mathbf{X}_{1}-\mathbf{X}_{2}\|_{\mathbb{H}}\qquad(\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}).$ (29)
The Ballion-Hadded theorem tells us that if $\ell$ is convex and $\beta$-smooth, then
$\langle\nabla\!\ell(\mathbf{X}_{1})-\nabla\!\ell(\mathbf{X}_{2}),\mathbf{X}_{1}-\mathbf{X}_{2}\ \rangle_{\mathbb{H}}\geqslant\frac{1}{\beta}\|\nabla\!\ell(\mathbf{X}_{1})-\nabla\!\ell(\mathbf{X}_{2})\|_{\mathbb{H}}^{2}\qquad(\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}).$
It is clear that (6) is convex, being the squared norm of a linear function. Moreover, comparing (11) and (26), we have
$\nabla\!\ell(\mathbf{X}_{1})-\nabla\!\ell(\mathbf{X}_{2})=\mathcal{K}(\mathbf{X}_{1})-\mathcal{K}(\mathbf{X}_{2})=\mathcal{K}(\mathbf{X}_{1}-\mathbf{X}_{2}).$ (30)
Therefore (30) becomes
$\|\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}\leqslant\beta\|\mathbf{X}\|_{\mathbb{H}}\qquad(\mathbf{X}\in\mathbb{H}),$ (31)
and (29) becomes
$\langle\mathcal{K}(\mathbf{X}),\mathbf{X}\ \rangle_{\mathbb{H}}\geqslant\frac{1}{\beta}\|\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}^{2}\qquad(\mathbf{X}\in\mathbb{H}).$ (32)
We next estimate $\beta$. It follows from (26) and the triangle inequality, that
$\|\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}\leqslant\|\mathbf{P}_{1}\mathbf{X}\|_{\mathbb{H}}+\|\mathbf{X}\mathbf{P}_{2}\|_{\mathbb{H}}\leqslant\big{(}\sigma_{\max}(\mathbf{P}_{1})+\sigma_{\max}(\mathbf{P}_{2})\big{)}\|\mathbf{X}\|_{\mathbb{H}}.$
Since $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are given by (27), we can take $\beta$ in (31) and (32) to be
$\beta=\sigma_{\max}(\mathbf{A}^{\top}\!\mathbf{A})+\lambda\sigma_{\max}((\mathbf{E}\mathbf{R})(\mathbf{E}\mathbf{R})^{\top})=\sigma_{\max}(\mathbf{A})^{2}+\lambda\sigma_{\max}(\mathbf{E}\mathbf{R})^{2}.$ (33)
We can now easily show that $\mathcal{G}$ is nonexpansive. Note that
$\|\mathcal{G}(\mathbf{X})\|_{\mathbb{H}}^{2}=\|\mathbf{X}-\gamma\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}^{2}=\|\mathbf{X}\|_{\mathbb{H}}^{2}+\gamma^{2}\|\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}^{2}-2\gamma\langle\mathbf{X},\mathcal{K}(\mathbf{X})\ \rangle_{\mathbb{H}}.$ (34)
Thus, if $\gamma$ is in the range given by (28), we have from (32) and (34) that
$\|\mathcal{G}(\mathbf{X})\|_{\mathbb{H}}^{2}\leqslant\|\mathbf{X}\|_{\mathbb{H}}^{2}-\gamma\left(\frac{2}{\beta}-\gamma\right)\|\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}^{2}\leqslant\|\mathbf{X}\|_{\mathbb{H}}^{2}.$ (35)
To complete the proof of Proposition (3), we must show that 1-1 is not an eigenvalue of $\mathcal{G}$. Indeed, if $1\in\sigma(\mathcal{G})$, we would have
$\mathcal{G}(\overline{\mathbf{X}})=\overline{\mathbf{X}}-\gamma\mathcal{K}(\overline{\mathbf{X}})=-\overline{\mathbf{X}},$
for some nonzero $\overline{\mathbf{X}}\in\mathbb{H}$. Thus, if $\gamma$ is in the range given by (28), we get from (31) that
$2\|\overline{\mathbf{X}}\|_{\mathbb{H}}=\gamma\|\mathcal{K}(\overline{\mathbf{X}})\|_{\mathbb{H}}<\frac{2}{\beta}\|\mathcal{K}(\overline{\mathbf{X}})\|_{\mathbb{H}}\leqslant 2\|\overline{\mathbf{X}}\|_{\mathbb{H}},$
However, this is not possible, implying that $1\notin\sigma(\mathcal{G})$.
Thus far, we focused on the gradient step operator, one of the two operators in (21). We next analyze the denoiser $\mathcal{D}$, which is defined in terms $\mathcal{W}$ and $\mathcal{V}$.
The properties of $\mathcal{W}$ and $\mathcal{V}$ are determined by that of $\mathbf{W}$ and $\mathbf{V}$; see (15) and (18). For instance, $\mathbf{W}$ and $\mathbf{V}$ are symmetric, and this forces $\mathcal{W}$ and $\mathcal{V}$ to be self-adjoint –– the choice of the inner product (4) is crucial in this regard. We will also require additional properties of $\mathcal{W}$ and $\mathcal{V}$ in our analysis, which are summarized in the following result.
Proposition 4.
The operators $\mathcal{V}$ and $\mathcal{W}$ are self-adjoint and nonexpansive, $\sigma(\mathcal{W})\subset[0,1]$, and $\mathrm{fix}\,(\mathcal{W})$ is a subspace of $\mathbb{H}$ consisting of matrices $\mathbf{F}$ of the form
$\mathbf{F}=\Big{[}c_{1}\boldsymbol{e}\ \cdots\ c_{L_{s}}\boldsymbol{e}\Big{]},$ (36)
where $c_{1},\ldots,c_{L_{s}}\in\mathbb{R}$ and $\boldsymbol{e}\in\mathbb{R}^{N_{m}}$ is the all-ones vector.
Proposition 4 is a key result that relies on the intricate construction of the denoisers in Section 3.3. Indeed, it is a non-trivial task to construct a kernel-based denoiser with the above mathematical properties. As previously discussed, the construction builds on prior works [42, 52], and the core part of the proof is based on Proposition 1. We have worked out the details in Appendix 7.1.
We can use the observation that $\mathcal{V}$ is nonexpansive to simplify our analysis. Recall the definition of $\mathcal{D}$ and $\mathcal{P}$ in (12) and (21). Using the associativity of composition, we can write
$\mathcal{P}=(\mathcal{V}\circ\mathcal{W})\circ\mathcal{G}=\mathcal{V}\circ(\mathcal{W}\circ\mathcal{G}).$
Therefore, if we can establish that $\mathcal{W}\circ\mathcal{G}$ is a contraction, it will follow that $\mathcal{P}$ is a contraction. In other words, we can focus on the two operators $\mathcal{W}$ and $\mathcal{G}$.
As a first observation, note that if we can find a nonzero $\overline{\mathbf{X}}$ that is a fixed point of both $\mathcal{W}$ and $\mathcal{G}$, then $\mathcal{W}\circ\mathcal{G}$ cannot be a contraction. Indeed, we would have $(\mathcal{W}\circ\mathcal{G})(\overline{\mathbf{X}})=\overline{\mathbf{X}}$, and consequently
$\|(\mathcal{W}\circ\mathcal{G})(\overline{\mathbf{X}})\|_{\mathbb{H}}=\|\overline{\mathbf{X}}\|_{\mathbb{H}},$
so that $\mathcal{W}\circ\mathcal{G}$ cannot be a contraction. In this connection, we have the following result.
Proposition 5.
There are no common nonzero fixed points of $\mathcal{W}$ and $\mathcal{G}$, namely, $\mathrm{fix}\,(\mathcal{W})\cap\mathrm{fix}\,(\mathcal{G})=\{\mathbf{0}\}$.
We know from Proposition 4 that the $\mathrm{fix}\,(\mathcal{W})$ consists of matrices $\mathbf{F}$ of the form (36). Thus, to establish the above result, it suffices to show that if $\mathbf{F}$ is of the form (36) and $\mathbf{F}\neq\mathbf{0}$, then $\mathbf{F}\notin\mathrm{fix}\,(\mathcal{G})$. Note that $\mathbf{F}\neq\mathbf{0}$ means that $c_{j}\neq 0$ for some $1\leqslant j\leqslant L_{s}$. We will show using contradiction that under this condition $\mathbf{F}\notin\mathrm{fix}\,(\mathcal{G})$. Suppose that $\mathcal{G}(\mathbf{F})=\mathbf{F}$. Then, from (26), we have $\mathcal{K}(\mathbf{F})=\mathbf{0}$, i.e.,
$\mathbf{P}_{1}\mathbf{F}+\mathbf{F}\mathbf{P}_{2}=\mathbf{0},$
where $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are given by (27). In particular, this means
$0=\langle\mathbf{F},\mathbf{P}_{1}\mathbf{F}+\mathbf{F}\mathbf{P}_{2}\rangle_{\mathbb{H}}=\langle\mathbf{F},\mathbf{P}_{1}\mathbf{F}\rangle_{\mathbb{H}}+\langle\mathbf{F}^{\top}\!,\mathbf{P}_{2}\mathbf{F}^{\top}\rangle_{\mathbb{H}}.$
Since $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are symmetric and positive semidefinite, we know that the two terms on the right are nonnegative. Therefore, we must have
$\langle\mathbf{F},\mathbf{P}_{1}\mathbf{F}\rangle_{\mathbb{H}}=0\quad\mbox{and}\quad\langle\mathbf{F}^{\top}\!,\mathbf{P}_{2}\mathbf{F}^{\top}\rangle_{\mathbb{H}}=0.$
We just need the first relation. Substituting (26) and noting that $\mathbf{P}_{1}=\mathbf{A}^{\top}\!\mathbf{A}$, we get
$0=\langle\mathbf{F},\mathbf{P}_{1}\mathbf{F}\rangle_{\mathbb{H}}=\mathrm{tr}(\mathbf{F}^{\top}\!\mathbf{P}_{1}\mathbf{F})=\big{(}c_{1}^{2}+\cdots+c_{L_{s}}^{2}\big{)}\|\mathbf{A}\boldsymbol{e}\|_{2}^{2},$
where $\|\cdot\|_{2}$ is the standard Euclidean norm on $\mathbb{R}^{N_{m}}$. In particular, since we assume that $c_{j}\neq 0$ for some $j$, we must have $\mathbf{A}\boldsymbol{e}=\mathbf{0}$. However, recall that $\mathbf{A}=\mathbf{S}\mathbf{B}$, where $\mathbf{S}$ and $\mathbf{B}$ are sampling and blur operators. Since $\mathbf{B}\boldsymbol{e}=\boldsymbol{e}$, i.e., blurring a constant-intensity image produces the same image, we get $\mathbf{0}=\mathbf{A}\boldsymbol{e}=\mathbf{S}(\mathbf{B}\boldsymbol{e})=\mathbf{S}\boldsymbol{e}$. This is the desired contradiction since we cannot get a null vector by sampling $\boldsymbol{e}$.
4.3 Main Result
It turns out that the condition $\mathrm{fix}\,(\mathcal{W})\cap\mathrm{fix}\,(\mathcal{G})=\{\mathbf{0}\}$ is not only necessary but also sufficient to guarantee that $\mathcal{W}\circ\mathcal{G}$ is a contraction, but under some assumptions on $\mathcal{W}$ and $\mathcal{G}$. To show this, we need the following result.
Lemma 1.
Suppose $\mathcal{T}\in\mathcal{L}(\mathbb{H})$ is self-adjoint and $\sigma(\mathcal{T})\subset(-1,1]$. Then, for any $\mathbf{X}\in\mathbb{H}$,
$\|\mathcal{T}(\mathbf{X})\|_{\mathbb{H}}=\|\mathbf{X}\|_{\mathbb{H}}\implies\mathcal{T}(\mathbf{X})=\mathbf{X}.$
In particular, if $\mathbf{X}\notin\mathrm{fix}\,(\mathcal{T})$, then $\|\mathcal{T}(\mathbf{X})\|_{\mathbb{H}}<\|\mathbf{X}\|_{\mathbb{H}}$.
The above result identifies a condition under which $\mathcal{T}$ acts as a contraction in the complement of the subspace $\mathrm{fix}\,(\mathcal{T})$. This is the closest an operator can come to being a contraction if it has a nontrivial $\mathrm{fix}\,(\mathcal{T})$, i.e., if $\mathrm{fix}\,(\mathcal{T})\neq\{\mathbf{0}\}$. Although the final conclusion in Lemma (1) appears in [49], we provide a proof in Appendix 7.2 for completeness. Importantly, we see from Propositions 3 and 4 that $\mathcal{W}$ and $\mathcal{G}$ satisfy the assumptions in Lemma 1.
We are now ready to prove our main result: the contractivity of $\mathrm{fix}\,(\mathcal{W})\cap\mathrm{fix}\,(\mathcal{G})$ and, consequently, the global linear convergence of HyDeFuse.
Theorem 1.
The fixed point operator $\mathcal{T}$ in (20) is a contraction. In particular, for any arbitrary initialization $\mathbf{X}_{0}$, the iterates of HyDeFuse generated by (9) converge linearly to the unique fixed point of $\mathcal{T}$.
We had earlier noted that $\mathcal{T}$ is a contraction if $\mathcal{W}\circ\mathcal{G}$ is a contraction. By Proposition 2, it suffices to show that for any $\mathbf{X}\neq\mathbf{0}$,
$\|(\mathcal{W}\circ\mathcal{G})(\mathbf{X})\|_{\mathbb{H}}<\|\mathbf{X}\|_{\mathbb{H}}.$
There are two possibilities. Supposex $\mathbf{X}\in\mathrm{fix}\,(\mathcal{G})$. Then $(\mathcal{W}\circ\mathcal{G})(\mathbf{X})=\mathcal{W}(\mathbf{X})$. However, we also have from Proposition 5, that $\mathbf{X}\notin\mathrm{fix}\,(\mathcal{W})$. Since $\mathcal{W}$ is self-adjoint and $\sigma(\mathcal{W})\subset(-1,1]$, we have from Lemma 1 that $\|\mathcal{W}(\mathbf{X})\|_{\mathbb{H}}<\|\mathbf{X}\|_{\mathbb{H}}$. Therefore,
$\|(\mathcal{W}\circ\mathcal{G})(\mathbf{X})\|_{\mathbb{H}}=\|\mathcal{W}(\mathbf{X})\|_{\mathbb{H}}<\|\mathbf{X}\|_{\mathbb{H}}.$
Consider the other possibility $\mathbf{X}\notin\mathrm{fix}\,(\mathcal{G})$. In this case, we have $\|\mathcal{G}(\mathbf{X})\|_{\mathbb{H}}<\|\mathbf{X}\|_{\mathbb{H}}$. Thus, since $\mathcal{W}$ is nonexpansive, we again have
$\|(\mathcal{W}\circ\mathcal{G})(\mathbf{X})\|_{\mathbb{H}}=\|\mathcal{W}(\mathcal{G}(\mathbf{X}))\|_{\mathbb{H}}\leqslant\|\mathcal{G}(\mathbf{X})\|_{\mathbb{H}}<\|\mathbf{X}\|_{\mathbb{H}}.$
Thus, we have established that $\mathcal{W}\circ\mathcal{G}$ and hence $\mathcal{T}$ is a contraction.
The second part of Theorem 1 is a standard application of the contraction mapping theorem [72], which guarantees the existence of a fixed point $\mathbf{X}^{*}\in\mathrm{fix}\,(\mathcal{T})$. In particular, if $\mu\in[0,1)$ is the contraction factor of $\mathcal{T}$, then
$\|\mathbf{X}_{k+1}-\mathbf{X}^{*}\|_{\mathbb{H}}=\|\mathcal{T}(\mathbf{X}_{k})-\mathcal{T}(\mathbf{X}^{*})\|_{\mathbb{H}}\leqslant\mu\|\mathbf{X}_{k}-\mathbf{X}^{*}\|_{\mathbb{H}}.$
In particular, we have
$\|\mathbf{X}_{k}-\mathbf{X}^{*}\|_{\mathbb{H}}\leqslant\mu^{k}\|\mathbf{X}_{0}-\mathbf{X}^{*}\|_{\mathbb{H}}\qquad(k\geqslant 0).$ (37)
This completes the proof of Theorem 1, establishing the global linear convergence of the iterates of HyDeFuse.
The above analysis allows us to identify the step size $\gamma$ in (9) required to guarantee convergence. This in turn depends on $\beta$ given (33). Since $\mathbf{A}=\mathbf{S}\mathbf{B}$, we can be shown that $\sigma_{\max}(\mathbf{A})\leqslant 1$, giving us the bound $\beta\leqslant 1+\lambda\sigma_{\max}(\mathbf{E}\mathbf{R})^{2}$, where $\sigma_{\max}(\mathbf{E}\mathbf{R})$ can easily be estimated using SVD. We can obtain a more precise bound using the power method [67] on the linear map $\mathcal{K}$ in (26). Empirical results show that this is often much smaller than the above bound. However, the analytical bound is useful as it is easy to compute and ensures convergence.
4.4 Discussion
After presenting our technical results, we now examine the connection between our analysis and related works [52, 42, 46, 49], which employ kernel denoisers to develop provably convergent PnP models. In [52, 42, 46], the kernel denoiser is designed as a proximal operator of a convex regularizer, allowing PnP to be linked to a convex optimization model. Standard results on proximal algorithms [72] are then used to establish objective convergence. However, in our case, the denoiser $\mathcal{D}$ is not necessarily self-adjoint, as $\mathcal{V}$ and $\mathcal{W}$ do not commute. This lack of symmetry implies that $\mathcal{D}$ cannot be expressed as the proximal operator of a convex regularizer, following a classical result by Moreau [73]. As a result, the approach in [52, 42, 46] is not directly applicable for HyDeFuse. Moreover, we establish a stronger result –– iterate convergence over objective convergence –– as well as the convergence rate. Our analysis instead builds on [49]. However, the loss function, the underlying vector space and operators, and the denoiser in HyDeFuse are different from the generic setup in [49]. As a result, the proofs from [49] do not directly apply, requiring us to refine and adapt the arguments.
5 Numerical Results
In this section, we validate our theoretical findings and demonstrate the empirical performance of HyDeFuse. Specifically, we provide numerical results on global convergence and the convergence rate, assess the proposed denoiser CasKD in comparison to existing denoisers, and benchmark HyDeFuse against various classical and state-of-the-art fusion methods.
5.1 Experimental Setup
We provide a detailed description of the dataset, the parameters of the forward model, the quality metrics used for evaluation, and the denoising and fusion methods used for comparison.
[…]
Blur models. The Starck-Murtagh filter was applied as the blurring operator for the HS image [8, 52]. The downsampling factors used to generate $\mathbf{Y}_{h}$ were set to 4 for Pavia, 6 for Chikusei, and 3 for Paris. The spectral response $\mathbf{R}$ was computed following the approach in [8]. If the SNR for HS and MS images is mentioned, it is denoted as SNRh and SNRm, respectively. If not specified, it is assumed to be 20 dB for both HS and MS across all datasets. We also compare HyDeFuse with recent methods using a Gaussian point spread function with radius $7$ and standard deviation $2$ [74]. This is to demonstrate that HyDeFuse performs effectively under varying blurring models.
Compared methods. We compare our kernel denoiser CasKD with both traditional and recent denoising algorithms. Specifically, we compare CasKD against BM4D [75], DnCNN [39], FHDD-NLM [52], KBR [51], LLRT [50], FastHyMix [54], and NGMeet [53]. We compare our fusion method HyDeFuse with a range of classical and state-of-the-art methods: GSA [7], CNMF [17], GLP [13], MAPSMM [9], HySure [8], GTTN [74], and CTDF [18]. Furthermore, within the PnP-PGD framework, we compare with denoisers such as DnCNN [39], BM4D [75], and GMM [45].
Quality metrics. We employed standard quality metrics [8, 43, 74], including Peak Signal-to-Noise Ratio (PSNR), Root Mean Squared Error (RMSE), Spectral Angle Mapper (SAM), Erreur Relative Globale Adimensionnelle de Synthèse (ERGAS), and Universal Image Quality Index (UIQI). For reconstruction tasks, lower RMSE, SAM, and ERGAS values indicate better performance, while higher PSNR and UIQI values (maximum is $1$) suggest superior reconstruction quality.
5.2 Denoising Results
Methods PSNR (0.0196) UIQI (0.0196) PSNR (0.0784) UIQI (0.0784)
Noisy 33.20 0.95 22.10 0.68
FHDD-NLM 43.22 0.985 35.45 0.97
BM4D 43.65 0.989 35.01 0.97
DnCNN 43.46 0.990 37.83 0.98
KBR 44.42 0.996 37.75 0.98
LLRT 43.07 0.994 $\mathbf{38.92}$ 0.97
FastHyMix 46.83 0.998 38.76 0.980
NGMeet $\mathbf{47.84}$ $\mathbf{0.998}$ $\mathbf{40.83}$ $\mathbf{0.983}$
CasKD $\mathbf{46.86}$ $\mathbf{0.998}$ 38.06 $\mathbf{0.987}$
Table 2: Denoising comparison on the Pavia dataset for Gaussian noise (standard deviation in brackets). The top two values are highlighted in bold, and the best value is underlined; this convention is followed in all the tables.
(a) ground truth
(b) FHDD-NLM
(c) BM4D
(d) DnCNN
(e) KBR
(f) LLRT
(g) FastHyMix
(h) NGMeet
(i) CasKD
Figure 4: Comparison of the proposed denoiser CasKD with existing denoisers on the Pavia dataset. The noise is Gaussian with standard deviation $0.019$. For visualization, only three bands are displayed.
Methods PSNR (0.0196) UIQI (0.0196) PSNR (0.0784) UIQI (0.0784)
Noisy 34.16 0.969 22.10 0.68
FHDD-NLM 44.77 0.977 35.52 0.969
BM4D 40.44 0.992 31.70 0.944
DnCNN 42.49 0.996 36.78 0.982
KBR 39.88 0.991 34.25 0.967
LLRT 41.08 0.997 35.82 0.981
FastHyMix $\mathbf{47.04}$ $\mathbf{0.998}$ 37.44 0.984
NGMeet $\mathbf{47.52}$ $\mathbf{0.998}$ $\mathbf{38.22}$ $\mathbf{0.986}$
CasKD 46.94 $\mathbf{0.998}$ $\mathbf{38.24}$ $\mathbf{0.987}$
Table 3: Same as Table 2 but for the Paris dataset.
The proposed denoiser CasKD is a combination of two kernel denoisers. Combining two denoisers does not always yield good results. However, we recall that the component denoisers act differently in CasKD –– they work on the inter- and intra-band correlations between patches. This is a possible explanation for the improved performance in Table 1. To further evaluate the performance of CasKD, we compare it with several existing denoisers: BM4D [75], DnCNN [39], FHDD-NLM [52], KBR [51], LLRT [50], FastHyMix [54], and NGMeet [53]. A visual comparison is shown in Figure 4. In Tables 2 and 3, we give a detailed comparison based on PSNR and UIQI values. We observe that CasKD performs competitively with these denoisers, and in some cases, it outperforms them.
5.3 Convergence Results
Figure 5: Convergence rate $\mu$ for different $\lambda$ appearing in the loss function (6). The parameters $\sigma_{1}$ and $\sigma_{2}$ are the standard deviations of the Gaussian (RBF) kernels used in $\mathcal{W}$ and $\mathcal{V}$.
We used the power method [67] to calculate the convergence rate of HyDeFuse. We see from (37) that the rate is given by $\mu$, the largest singular value of the linear operator $\mathcal{P}$. However, computing $\mu$ via SVD is impractical because $\mathcal{P}$ is too large to store. Instead, the power method estimates $\mu$ by iteratively applying $\mathcal{P}$ and its adjoint $\mathcal{P}^{*}$ [67], where it follows from (21) that $\mathcal{P}^{*}=\mathcal{G}\circ\mathcal{W}\circ\mathcal{V}$, since $\mathcal{V},\mathcal{W}$ and $\mathcal{G}$ are self-adjoint.
In Table 4, we use the Chikusei dataset to validate the contraction factor of $\mathcal{P}$ in HyDeFuse. This depends on the step size $\gamma$, the forward model parameters $\mathbf{A}$ and $\mathbf{R}$, and the denoiser $\mathcal{D}$. Specifically, we used the power method to compute $\mu$ for step sizes $0<\gamma<2/\beta$ and for different denoiser settings. We see that as predicted by Theorem 1, the contraction factor is <1<1. On the same dataset, we computed the contraction factor for different values of $\lambda$, which is used to balance the loss function. The results are shown in Figure 5. Again, we see that $\mu$ always remains below $1$. Additionally, we see that convergence is faster when $\lambda$ is neither too large nor too small, which is noteworthy as this range is also expected to yield optimal fusion performance.
Figure 6: The plot shows the successive differences between HyDeFuse iterates and the corresponding PSNR values for different step sizes $\gamma$ (expressed in units of $1/\beta$). We computed $\beta$ using the power method and found it to be $0.5998$ for this setup. We observe that convergence is achieved for $\gamma\in\left(0,2/\beta\right)$. However, as demonstrated with the example $\gamma=2.2/\beta$, convergence cannot be guaranteed when $\gamma$ falls outside this range. Note that the PSNR stabilizes within 10 iterations for optimally chosen $\gamma$.
Figure 7: Same as Figure 6 but for the Pavia dataset and using a Gaussian point spread function with radius $7$ and a standard deviation of 22 for the blur operator $\mathbf{B}$. In this case, $\beta=0.389$. Once again, empirical convergence is observed for $\gamma\in\left(0,2/\beta\right)$, while values of $\gamma$ outside this range lead to divergence. Notice that the PSNR stabilizes within 10 iterations for an optimal $\gamma$.
(a) ground truth
(b) $\mathbf{X}_{0}$ : all-ones
(c) $\mathbf{X}_{0}$ : all-zeros
(d) $\mathbf{X}_{0}$ : white noise
Figure 8: Fusion results on the Chikusei dataset. The reconstructions shown are obtained from HyDeFuse using different initializations $\mathbf{X}_{0}$. In all cases, the performance metrics remain identical: PSNR = 42.27, RMSE = 0.0076, ERGAS = 1.39, SAM = 1.79, and UIQI = 0.967.
$\gamma$ $(a,a)$ $(b,a)$ $(c,a)$ $(a,b)$ $(b,b)$ $(c,c)$
0.1 0.985 0.986 0.986 0.986 0.986 0.986
0.5 0.961 0.964 0.964 0.964 0.964 0.964
1 0.922 0.929 0.929 0.929 0.929 0.929
1.5 0.884 0.894 0.894 0.894 0.894 0.894
1.8 0.861 0.873 0.873 0.873 0.873 0.873
1.9 0.906 0.897 0.897 0.897 0.897 0.896
Table 4: Contraction factors for different step sizes and denoiser settings. The step size $\gamma$ is given in multiples of $1/\beta$; $\beta$ is computed using the power method and found to be $0.9757$. The denoiser settings $(\sigma_{1},\sigma_{2})$ represent the standard deviations of the Gaussian (RBF) kernels in $\mathcal{W}$ and $\mathcal{V}$, where $a=0.0196$, $b=0.039$, and $c=0.117$ on a scale of $[0,1]$.
As further evidence of the contractive nature of the reconstruction operator $\mathcal{T}$, we conducted fusion experiments on the Chikusei dataset. The results show that, irrespective of the initialization, the iterations consistently converge to the same fixed point, confirming the uniqueness of the fixed points of the contractive operator $\mathcal{T}$.
Finally, we analyze the convergence of the iterates $\{\mathbf{X}_{k}\}$ generated by (9). A sample result on the Paris dataset for the Starck-Murtagh filter as the blurring operator is presented in Figure 6. A similar result is reported in Figure 7, but for Pavia and using a Gaussian point spread function as the blur model. As predicted by Theorem 1, the successive difference $\|\mathbf{X}_{k}-\mathbf{X}_{k-1}\|$ approaches zero, which is confirmed by these figures. Importantly, observe that the PSNR stabilizes within 10–15 iterations. This pattern was consistently observed across all three datasets. An interesting finding is that when the step size $\gamma$ exceeds the admissible range in (28), the algorithm diverges, causing $\|\mathbf{X}_{k}-\mathbf{X}_{k-1}\|$ to increase gradually. This highlights the importance of maintaining an appropriate step size. Moreover, Figure 6 suggests that increasing $\gamma$ does not necessarily accelerate convergence. In fact, notice that the term $\gamma(2/\beta-\gamma)$ in (35) governs the nonexpansivity of $\mathcal{G}$; this reaches its maximum when $\gamma$ is neither too small nor too large.
5.4 Fusion Results
A comparison between HyDeFuse and various classical methods, including GSA [7], CNMF [17], GLP [13], MAPSMM [9], and HySure [8] is provided in Figure 9. The corresponding metrics are reported in Table 7. It is not surprising that HyDeFuse outperforms these classical methods.
We next compare HyDeFuse with two recent methods, GTTN [74] and CTDF [18], on the Pavia dataset. A visual comparison is provided in Figure 10, while various performance metrics are presented in Tables 5 and 6. We observe that our method, HyDeFuse, performs on par with these approaches. To further demonstrate its effectiveness and generalizability, we evaluate HyDeFuse under two different SNR settings—one used in the original papers of GTTN and CTDF, and another that we have consistently used for comparisons.
(a) ground truth
(b) Bicubic
(c) CNMF
(d) GLP
(e) MASMM
(f) GSA
(g) Hysure
(h) HyDeFuse
Figure 9: Fusion results on the Pavia dataset. See Table 7 for the quality metrics.
(a) ground truth
(b) GTNN
(c) CTDF
(d) HyDeFuse
Figure 10: Comparison of recent tensor-based methods with Gaussian point spread function (model for $\mathbf{B}$) with radius $7$ and standard deviation 22. The corresponding metrics are shown in Table 5.
Method PSNR RMSE ERGAS SAM UIQI
CTDF 38.49 0.013 1.94 3.67 0.9869
GTNN $\mathbf{41.58}$ $\mathbf{0.008}$ $\mathbf{1.37}$ $\mathbf{2.39}$ $\mathbf{0.9859}$
HyDeFuse $\mathbf{40.79}$ $\mathbf{0.009}$ $\mathbf{1.33}$ $\mathbf{2.03}$ $\mathbf{0.994}$
Table 5: Quality metrics for Figure 10. In this experiment, $\text{SNR}_{h}=30\,\text{dB}$ and $\text{SNR}_{m}=35\,\text{dB}$.
Method PSNR RMSE ERGAS SAM UIQI
CTDF 28.37 0.038 6.28 11.82 0.88
GTNN $\mathbf{31.53}$ $\mathbf{0.027}$ $\mathbf{4.60}$ $\mathbf{8.15}$ $\mathbf{0.9325}$
HyDeFuse $\mathbf{34.48}$ $\mathbf{0.018}$ $\mathbf{2.92}$ $\mathbf{4.17}$ $\mathbf{0.972}$
Table 6: Quality metrics with Gaussian point spread function (model for $\mathbf{B}$) with radius $7$ and standard deviation 22. In this experiment, $\text{SNR}_{h}=20\,\text{dB}$ and $\text{SNR}_{m}=20\,\text{dB}$.
Dataset Methods PSNR RMSE ERGAS SAM UIQI
Bicubic 25.05 0.0559 8.44 8.40 0.746
MASMM 28.63 0.0384 5.85 9.61 0.91
GLP 28.39 0.037 5.95 9.83 0.9018
Pavia CNMF 29.46 0.033 5.28 8.09 0.924
GSA 29.95 0.0318 5.03 9.98 0.918
HySure $\mathbf{34.12}$ $\mathbf{0.020}$ $\mathbf{3.029}$ $\mathbf{5.398}$ $\mathbf{0.965}$
HyDeFuse $\mathbf{35.89}$ $\mathbf{0.016}$ $\mathbf{2.39}$ $\mathbf{3.61}$ $\mathbf{0.980}$
Bicubic 23.05 0.0702 6.24 6.21 0.57
MASMM 25.01 0.0556 5.06 5.59 0.7761
GLP 24.36 0.059 5.43 6.35 0.751
Paris CNMF 25.39 0.05 4.85 4.55 0.80
GSA 24.48 0.06 5.36 5.78 0.78
HySure $\mathbf{26.82}$ $\mathbf{0.04}$ $\mathbf{4.13}$ $\mathbf{3.17}$ $\mathbf{0.839}$
HyDeFuse $\mathbf{27.24}$ $\mathbf{0.04}$ $\mathbf{3.93}$ $\mathbf{2.81}$ $\mathbf{0.846}$
Bicubic 29.02 0.035 7.49 7.30 0.54
MASMM 30.84 0.028 6.51 8.74 0.7170
GLP 30.20 0.030 6.80 8.31 0.698
Chikusei CNMF 34.14 0.019 3.97 5.23 0.847
GSA 32.29 0.020 6.89 8.84 0.761
HySure $\mathbf{40.56}$ $\mathbf{0.009}$ $\mathbf{1.99}$ $\mathbf{2.58}$ $\mathbf{0.937}$
HyDeFuse $\mathbf{42.27}$ $\mathbf{0.007}$ $\mathbf{1.39}$ $\mathbf{1.79}$ $\mathbf{0.967}$
Table 7: Comparison with classical methods for all three datasets.
(a) ground truth
(b) DnCNN
(c) BM4D
(d) GMM
(e) CasKD
Figure 11: PnP-PGD fusion on the Chikusei dataset where we compare our denoiser CasKD with other denoisers (as regularizers). The metrics are reported in Table 8.
Denoisers PSNR RMSE ERGAS SAM UIQI
DnCNN 42.29 0.0076 1.43 1.90 0.961
BM4D 42.40 0.0075 1.47 1.97 0.967
GMM 35.67 0.016 2.78 5.25 0.872
CasKD 42.27 0.008 1.39 1.79 0.967
Table 8: Quality metrics for Figure 11.
Finally, we analyze the impact of replacing our CasKD denoiser in HyDeFuse with existing denoisers, including DnCNN [39], BM4D [75], and GMM [45]. As shown in Figure 11, the reconstruction quality of CasKD is comparable to that of DnCNN and BM4D, and in some cases, it even surpasses them. The reported fused images were obtained after 100 iterations. We note that it is not expected that a simple kernel denoiser such as CasKD –– that essentially works by averaging local pixels –– can beat a trained denoiser such as DnCNN or a sophisticated denoiser such as BM4D. However, note that we are using CasKD as a regularizer not as a denoiser, where it is applied repeatedly across the iterations. This iterative regularization is what enables CasKD to achieve strong fusion results.
Dataset SNR Denoiser PSNR RMSE ERGAS SAM UIQI
$\mathcal{W}$ 33.33 0.021 3.04 5.19 0.960
Pavia 20dB CasKD 35.89 0.016 2.39 3.61 0.980
$\mathcal{V}$ 35.36 0.017 2.60 3.88 0.977
$\mathcal{W}$ 38.81 0.011 2.09 3.19 0.927
Chikusei 20dB CasKD 42.27 0.0076 1.39 1.79 0.967
$\mathcal{V}$ 42.17 0.0078 1.42 1.84 0.963
Table 9: Comparison of CasKD and its component denoisers as implicit regularizers in HyDeFuse.
6 Conclusion
The objective of this work was to propose an iterative denoising-based algorithm for hyperspectral fusion––modeled on the plug-and-play framework––that not only exhibits strong empirical performance but also comes with a provable convergence guarantee. Specifically, we utilized kernel denoisers, which provide effective noise reduction while maintaining tractable mathematical properties. We showed that there is scope for innovation within this class and came up with a cascaded denoiser, CasKD, which outperforms all previously reported kernel denoisers for hyperspectral images. Moreover, we demonstrated that the performance of CasKD is comparable with existing state-of-the-art (non-trained) denoisers. As explained in the introduction, we do not consider trained deep denoisers since establishing predictable mathematical properties and coming up with convergence guarantees is fundamentally challenging.
To integrate the implicit CasKD regularizer with a model-based loss for hyperspectral fusion, we employed the classical proximal gradient descent framework. Notably, leveraging the mathematical properties of CasKD, we rigorously and unconditionally established global linear convergence of our iterative fusion algorithm, HyDeFuse. We empirically validated the theoretical guarantee and analyzed how various parameters of HyDeFuse influence the convergence rate. While iterate convergence ensures the stability and reliability of the algorithm, it does not inherently guarantee high-quality reconstruction. We conducted extensive experiments on multiple hyperspectral datasets, demonstrating that HyDeFuse outperforms classical methods and performs competitively with recent state-of-the-art approaches. In particular, we observed that for proper settings of the internal parameters, the PSNR of the reconstruction stabilizes within 10–15 iterations. Although this was not discussed in the present work, the algorithmic concepts and mathematical guarantees can be extended to other models for denoising-driven regularization. It remains to be seen whether they can match state-of-the-art methods.
7 Appendix
7.1 Proof of Proposition 4
We recall that $\mathcal{V}$ is the bandwise denoiser given by (15) and $\mathcal{W}$ is the high-dimensional denoiser given by (18). The difference is that a difference $\mathbf{W}^{(b)}$ is used for each band in $\mathcal{V}$, while the same $\mathbf{W}$ is used across all bands in $\mathcal{W}$. To show that $\mathcal{V}$ is self-adjoint, we need to check that, for all $\mathbf{X},\mathbf{Y}\in\mathbb{H}$,
$\big{\langle}\mathcal{V}(\mathbf{X}),\mathbf{Y}\big{\rangle}_{\mathbb{H}}=\big{\langle}\mathbf{X},\mathcal{V}(\mathbf{Y})\big{\rangle}_{\mathbb{H}}.$
Recall that $\mathbb{H}$ is the space of matrices of size $N_{m}\times L_{s}$. In this part, it will be convenient to use $\boldsymbol{x}_{b}$ and $\boldsymbol{y}_{b}$ for the $b$-th columns (vectors in $\mathbb{R}^{N_{m}}$) of $\mathbf{X}$ and $\mathbf{Y}$. Using this notation, we can write (4) as
$\langle\mathbf{X},\mathbf{Y}\rangle_{\mathbb{H}}=\sum_{b=1}^{L_{s}}\boldsymbol{x}_{b}^{\top}\boldsymbol{y}_{b}.$
As the $b$-th band of $\mathcal{V}(\mathbf{X})$ is $\mathbf{W}^{(b)}\boldsymbol{x}_{b}$ and $\mathbf{W}^{(b)}$ is symmetric (Proposition 1), we have
$\big{\langle}\mathcal{V}(\mathbf{X}),\mathbf{Y}\big{\rangle}_{\mathbb{H}}=\sum_{b=1}^{L_{s}}(\mathbf{W}^{(b)}\boldsymbol{x}_{b})^{\top}\boldsymbol{y}_{b}=\sum_{b=1}^{L_{s}}\boldsymbol{x}_{b}^{\top}(\mathbf{W}^{(b)}\boldsymbol{y}_{b})=\big{\langle}\mathbf{X},\mathcal{V}(\mathbf{Y})\big{\rangle}_{\mathbb{H}}.$
The verification for $\mathcal{W}$ is identical. Thus, $\mathcal{V}$ and $\mathcal{W}$ are self-adjoint.
That $\mathcal{V}$ and $\mathcal{W}$ are nonexpansive follows directly from the nonexpansive property of $\mathbf{W}^{(b)}$ and $\mathbf{W}$. Indeed, since these matrices are stochastic and symmetric, it is not difficult to show that they are nonexpansive. Hence, for any $\mathbf{X}\in\mathbb{H}$, we have
$\|\mathcal{V}(\mathbf{X})\|^{2}_{\mathbb{H}}=\sum_{b=1}^{L_{s}}\|\mathbf{W}^{(b)}\boldsymbol{x}_{b}\|^{2}_{2}\leqslant\sum_{b=1}^{L_{s}}\|\boldsymbol{x}_{b}\|^{2}_{2}=\|\mathbf{X}\|^{2}_{\mathbb{H}}.$
where $\|\cdot\|_{2}$ is the Euclidean norm on $\mathbb{R}^{N_{m}}$. The proof for $\mathcal{W}$ is identical.
We know from the spectral theory of linear operators that a self-adjoint operator has real eigenvalues [67]. Moreover, since we have shown that $\mathcal{W}$ is nonexpansive, we can conclude that if $\lambda\in\sigma(\mathcal{W})$, then $|\lambda|\leqslant 1$. In other words, $\sigma(\mathcal{W})\subset[-1,1]$. We know from Proposition 1 that $\mathbf{W}$ is symmetric positive semidefinite. Therefore, for any $\mathbf{X}\in\mathbb{H}$,
$\langle\mathcal{W}(\mathbf{X}),\mathbf{X}\rangle_{\mathbb{H}}=\sum_{b=1}^{L_{s}}(\mathbf{W}\boldsymbol{x}_{b})^{\top}\boldsymbol{x}_{b}=\sum_{b=1}^{L_{s}}\boldsymbol{x}_{b}^{\top}\mathbf{W}\boldsymbol{x}_{b}\geqslant 0.$
This means that every eigenvalue of $\mathcal{W}$ is nonnegative. Hence $\sigma(\mathcal{W})\subset[0,1]$.
Finally, we will show that if $\mathcal{W}(\mathbf{X})=\mathbf{X}$, then each band $\boldsymbol{x}_{b}$ must be a multiple of $\boldsymbol{e}$, the all-ones vector in $\mathbb{R}^{N_{m}}$. This follows directly from the fact that $1$ is a simple eigenvalue of $\mathbf{W}$. Indeed, since $\mathbf{W}$ is stochastic and irreducible (Proposition 1), the Perron-Frobenius theorem tells us that $1$ must be a simple eigenvalue of $\mathbf{W}$ [67]. Now, note that the equation $\mathcal{W}(\mathbf{X})=\mathbf{X}$ translates to $\mathbf{W}\boldsymbol{x}_{b}=\boldsymbol{x}_{b}$ for all $b$, i.e., $\boldsymbol{x}_{b}=\mathbf{0}$ or $\boldsymbol{x}_{b}$ is an eigenvector with eigenvalue $1$. In the latter case, $\boldsymbol{x}_{b}=c_{b}\boldsymbol{e}$ for some $c_{b}\in\mathbb{R}$. This establishes (36), also completing the proof of Proposition 4.
7.2 Proof of Lemma 1
Since $\mathcal{T}$ is a self-adjoint operator on the inner product space $\mathbb{H}$, we can diagonalize it on an orthonormal basis [67]. More specifically, if $d$ is the dimension of $\mathbb{H}$, then we can find $\lambda_{1},\ldots,\lambda_{d}\in\mathbb{R}$ and an orthonormal basis $\mathbf{H}_{1},\ldots,\mathbf{H}_{d}$ of $\mathbb{H}$ such that, for any $\mathbf{X}\in\mathbb{H}$,
$\mathbf{X}=\sum_{j=1}^{d}\langle\mathbf{X},\mathbf{H}_{j}\rangle_{\mathbb{H}}\ \mathbf{H}_{j}\quad\mbox{and}\quad\mathcal{T}(\mathbf{X})=\sum_{j=1}^{d}\lambda_{j}\langle\mathbf{X},\mathbf{H}_{j}\rangle_{\mathbb{H}}\ \mathbf{H}_{j}.$ (38)
In particular,
$\|\mathbf{X}\|^{2}_{\mathbb{H}}=\sum_{j=1}^{d}\langle\mathbf{X},\mathbf{H}_{j}\rangle_{\mathbb{H}}^{2}\quad\mbox{and}\quad\|\mathcal{T}(\mathbf{X})\|^{2}_{\mathbb{H}}=\sum_{j=1}^{d}\lambda^{2}_{j}\langle\mathbf{X},\mathbf{H}_{j}\rangle_{\mathbb{H}}^{2}.$ (39)
Let $\mathbf{X}\in\mathbb{H}$. We have from (39) that $\|\mathcal{T}(\mathbf{X})\|_{\mathbb{H}}=\|\mathbf{X}\|_{\mathbb{H}}$ if and only if
$\sum_{j=1}^{d}(1-\lambda^{2}_{j})\langle\mathbf{X},\mathbf{H}_{j}\rangle_{\mathbb{H}}^{2}=0.$
However, $\lambda_{j}\in(-1,1]$ by assumption. Therefore, we must have $\lambda_{j}=1$ or $\langle\mathbf{X},\mathbf{H}_{j}\rangle_{\mathbb{H}}=0$. In either case, it follows from (38) that $\mathcal{T}(\mathbf{X})=\mathbf{X}$. This completes the proof of Lemma 1.